# Frostman random variables, entropy inequalities, and applications

### v3

@@ 4.1 Diagonal forms with independent random variables
$\begin{split}&\max\{H_{n}(X+Y),~H_{n}(\alpha X^{2}+\beta Y^{2})\}\geq n(s+\epsilon_{1})\\ &\qquad\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).\end{split}$ (39)
for all $n\in\mathbb{N}$.
@@
Write $\phi(x,y)=p(x)+q(y)$, where $p(x)=\alpha x^{2}$ and $q(y)=\beta y^{2}$ are functions whose domains are both taken to be $c_{2},-c_{1}]\cup[c_{2},c_{1}]$. One asserts by Example 2.16 that
$\begin{split}H_{n}(p(X))&=H_{n}(X)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1),\\ H_{n}(q(Y))&=H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).\end{split}$ (40)
Define $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y),$ with $|\alpha|X$ and $|\beta|Y$ being $(s,O_{\phi}(C_{1}))$-Frostman and $(s,O_{\phi}(C_{2}))$-Frostman, respectively. Let $W_{1}=(\sqrt{|\alpha|}X_{1},\sqrt{|\beta|}Y_{1})$, $W_{2}=(\sqrt{|\alpha|}X_{2},\sqrt{|\beta|}Y_{2})$ be i.i.d. copies of $W$. Then $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$, and $X_{1},X_{2},Y_{1},Y_{2}$ are independent.
@@
From Example 2.11 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O_{\phi}(c_{1})$-Lipschitz. We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O_{\phi}(C_{1})$-Lipschitz function
@@
$\leq H_{n}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
$\qquad+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1)$
$\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1).$ (41)
We now estimate the terms in the last expression. For the first term, apply Lemma 2.21 to get that
$H_{n}(X_{1}+X_{2})$ $\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1),$
$=2H_{n}(X+Y)-H_{n}(Y)+O(1).$
One can argue similarly for $H_{n}(Y_{1}+Y_{2})$, hence,
$H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})\leq 4H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (42)
For the third term, Lemma 2.29 shows that
$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))\leq 3H_{n}(p(X)+q(Y))$
$\qquad\qquad\qquad-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$ (43)
For the fourth term, applying Lemma 2.21 shows that
$H_{n}(X_{1}-X_{2})\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1).$
So similarly
$H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\leq 4H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (44)
@@
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+$ $5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}\alpha^{-s_{1}}\beta^{-s_{2}})$-Frostman. Here $\log(\alpha^{-s_{1}}\beta^{-s_{2}})=O_{\phi}(1)$. Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O_{\phi,c_{1}}(1)$. So we can deduce by Lemma 2.32 and Theorem 2.41 that
$H_{n}\big{(}X\big{)}\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}\big{(}Y\big{)}\geq ns_{2}-\log C_{2}+O(1),$
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}\geq n(s+\epsilon_{0})+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O_{\phi,s}(1)$
@@
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n\left(s+\frac{\epsilon_{0}}{11}\right)$
$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
Hence we complete the proof of the lemma with constant $\epsilon_{1}=\epsilon_{0}/11$.
@@
$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))\leq 9H_{n}(p(X)+q(Y))$
$\qquad\qquad\qquad-4H_{n}(X)-4H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
@@
We emphasize that the parameters $C_{1},C_{2}$ in Lemma 4.1 can be unbounded as $n$ tends to infinity. And, when we apply the above lemma in the next subsection, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms. So it cannot be replaced by $O_{C_{1},C_{2}}(1)$.
4.2 Diagonal forms with dependent random variables
Now we prove Theorem 1.2 in the following form, where $X$ and $Y$ are allowed to be dependent. We still take $\phi(x,y)=\alpha x^{2}+\beta y^{2}$ in this subsection. And a constant depending on $\phi$ means it depends on $\alpha$ and $\beta$.
Theorem 1.2’. Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C,c>0$. There is a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1).$
@@ 4.2 Diagonal forms with dependent random variables
For simplicity, we write
$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}-ns,$
and our aim is to prove that $\Psi\geq\epsilon n$ for some positive $\epsilon$. In particular,
$\frac{H_{n}(X)+H_{n}(Y)}{2},\,\,H_{n}(X+Y),\,\,H_{n}(\alpha X^{2}+\beta Y^{2})\,\leq\,ns+\Psi.$ (45)
By Lemma 2.30, the random variables $X$, $Y$, $(X,Y)$ are $(s_{1},C\cdot\max\{1,c\})$-Frostman, $(s_{2},C\cdot\max\{1,c\})$-Frostman and $(s_{1}+s_{2},C)$-Frostman, respectively.

### v4

@@ 4.1 Diagonal quadratics with independent random variables
$H_{n}(X+Y)+H_{n}(\alpha X^{2}+\beta Y^{2})\geq n(2s+\epsilon_{1})+O(\log\mathfrak{K})$
for all $n\in\mathbb{N}$, where
$\log\mathfrak{K}=O\big(\log_{+}C_{1}C_{2}+\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\alpha|\big|+\big|\log|\beta|\big|\big)+O_{s}(1)$
@@
Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$, where $\rho_{1}(x)=\alpha x^{2}$ and $\rho_{2}(y)=\beta y^{2}$. One asserts by Example 2.12 and Lemma 2.16 that
$\begin{split}H_{n}(\rho_{1}(X))&=H_{n}(X)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\alpha|\big|+1\big)=O(\log\mathfrak{K}),\\ H_{n}(\rho_{2}(Y))&=H_{n}(Y)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\beta|\big|+1\big)=O(\log\mathfrak{K}).\end{split}$ (36)
Define $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y)$. Let $W_{1}=(\sqrt{|\alpha|}X_{1},\sqrt{|\beta|}Y_{1})$, $W_{2}=(\sqrt{|\alpha|}X_{2},\sqrt{|\beta|}Y_{2})$ be i.i.d. copies of $W$. Then $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$, and $X_{1},X_{2},Y_{1},Y_{2}$ are independent.
@@
From Example 2.11 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O(|\alpha|c_{1}+|\beta|c_{1}+1)$-Lipschitz. We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O(|\alpha|c_{1}+|\beta|c_{1})$-Lipschitz function
@@
$\leq H_{n}(X_{1}+X_{2},\,Y_{1}+Y_{2},\,\rho_{1}(X_{1})+\rho_{1}(X_{2})+\rho_{2}(Y_{1})+\rho_{2}(Y_{2}))$
$\qquad+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O(\log\mathfrak{K})$
$\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})$
$\qquad+H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})+\rho_{2}(Y_{1})+\rho_{2}(Y_{2}))+O(\log\mathfrak{K}),$ (37)
We now estimate the terms in the last expression. Lemma 2.21 shows that
$H_{n}(X_{1}\pm X_{2})$ $\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1),$
$=2H_{n}(X+Y)-H_{n}(Y)+O(1),$ (38)
and similarly
$H_{n}(Y_{1}\pm Y_{2})\leq 2H_{n}(X+Y)-H_{n}(X)+O(1).$ (39)
And Lemma 2.29 tells that
$H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})+\rho_{2}(Y_{1})+\rho_{2}(Y_{2}))$
$\qquad\leq 3H_{n}(\rho_{1}(X)+\rho_{2}(Y))-H_{n}(X)-H_{n}(Y)+O(\log\mathfrak{K}).$ (40)
@@
$H_{n}\big(|W_{1}-W_{2}|^{2}\big)+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$
Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})$-Frostman. Here $\log(|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})=O(\log\mathfrak{K})$ since $s_{1},s_{2}$ are bounded. Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O\big(\sqrt{|\alpha|}c_{1}+\sqrt{|\beta|}c_{1}\big)$. So we can deduce by Lemma 2.32 and Theorem 2.41 that
$\begin{split}&H_{n}\big(X\big)\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}\big(Y\big)\geq ns_{2}-\log C_{2}+O(1),\\ &\qquad\qquad\qquad H_{n}\big(|W_{1}-W_{2}|^{2}\big)\geq n(s+\epsilon_{0})+O(\log\mathfrak{K})\end{split}$ (41)
@@
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n\left(s+\frac{\epsilon_{0}}{11}\right)+O(\log\mathfrak{K}).$ (42)
On the other hand, we can also deduce by independence, (9), (36) and (41) that
$\min\{H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\}$
$\qquad\geq\min\Big\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,\frac{H_{n}(\alpha X^{2})+H_{n}(\beta Y^{2})}{2}\Big\}+O(1)$
$\qquad\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+O(\log\mathfrak{K})\geq ns+O(\log\mathfrak{K}).$ (43)
Putting (42) and (43) together, we complete the proof of the lemma with constant $\epsilon_{1}=\epsilon_{0}/11$.
@@
$H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})-\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))\leq 9H_{n}(\rho_{1}(X)+\rho_{2}(Y))$
$\qquad\qquad\qquad-4H_{n}(X)-4H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K}).$
@@
We emphasize that $C_{1},C_{2}$ in Lemma 4.1 can be unbounded as $n$ tends to infinity. And, when we apply the above lemma in the following subsections, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms. So it cannot be replaced by $O_{C_{1},C_{2}}(1)$.
Lemma 4.3.
Let
$\rho_{1}(x)=\alpha_{2}x^{2}+\alpha_{1}x+\alpha_{0},\quad\rho_{2}(y)=\beta_{2}x^{2}+\beta_{1}x+\beta_{0},$
where $\alpha_{2},\beta_{2}\in{\mathbb{R}}^{\ast}$ and $\alpha_{1},\alpha_{0},\beta_{1},\beta_{0}\in{\mathbb{R}}$. Let $1/2<s<1$ and $C_{1},C_{2},c>0$. There exists a constant $\epsilon_{1}^{\prime}=\epsilon_{1}^{\prime}(s)>0$ such that the following statement holds: Let $X$, $Y$ be independent random variables with ranges contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(2s+\epsilon_{1}^{\prime})+O(\log\mathfrak{K})$
for all $n\in\mathbb{N}$, where
$\log\mathfrak{K}$ $=O_{s}\big(\log_{+}C_{1}+\log_{+}C_{2}+1\big)$
$+O\big(|\log|\alpha_{2}|\big|+\big|\log|\beta_{2}|\big|+\log_{+}|\alpha_{1}|+\log_{+}|\beta_{1}|+\log_{+}c).$
@@
Note that by completing the squares, we have
$\rho_{1}(x)=\alpha_{2}\left(x+a\right)^{2}+\alpha_{0}-\frac{\alpha_{1}^{2}}{4\alpha_{2}},\quad\rho_{2}(y)=\beta_{2}\left(x+b\right)^{2}+\beta_{0}-\frac{\beta_{1}^{2}}{4\beta_{2}},$
where $a=\frac{\alpha_{1}}{2\alpha_{2}}$ and $b=\frac{\beta_{1}}{2\beta_{2}}$. By Lemma 2.16,
$\begin{split}H_{n}(X+Y)&=H_{n}((X+a)+(Y+b))+O(1),\\ H_{n}(\rho_{1}(X)+\rho_{2}(Y))&=H_{n}(\alpha_{2}(X+a)^{2}+\beta_{2}(Y+b)^{2})+O(1).\end{split}$ (44)